# Generalized $T\bar{T}$-like flows for scalar theories in two dimensions

### v1

@@ Appendix A Details of the Lagrangian, and the Energy-Momentum Tensor
×$\times\Bigl{(}3\mathcal{P}^{2}-13P_{1}^{2}+\cosh(4\gamma)(\mathcal{P}^{2}-3P_{1}^{2})+4\cosh(2\gamma)(\mathcal{P}^{2}+4P_{1}^{2})-16\bigl{(}\cosh(\gamma)\bigr{)}^{3}\mathcal{P}P_{1}\sinh(\gamma)\Bigr{)}$
$+\frac{4d_{4}^{4}}{\mathcal{P}}\biggl{(}6\mathrm{csch}(\gamma)(\mathcal{P}-P_{1})P_{1}^{2}(\mathcal{P}+P_{1})-5\bigl{(}\mathrm{csch}(\gamma)\bigr{)}^{7}\mathcal{P}^{4}-\bigl{(}\mathrm{csch}(\gamma)\bigr{)}^{3}P_{1}^{2}(P_{1}^{2}-6\mathcal{P}^{2})+\bigl{(}\mathrm{csch}(\gamma)\bigr{)}^{5}(6\mathcal{P}^{2}P_{1}^{2}-4\mathcal{P}^{4})$
$+P_{1}^{2}\mathrm{sech}(\gamma)\Bigl{(}6(P_{1}^{2}-\mathcal{P}^{2})\tanh(\gamma)+(7P_{1}^{2}-12\mathcal{P}^{2})\bigl{(}\mathrm{sech}(\gamma)\bigr{)}^{2}\tanh(\gamma)+4P_{1}\bigl{(}\mathrm{sech}(\gamma)\bigr{)}^{4}\bigl{(}8\mathcal{P}+5P_{1}\tanh(\gamma)\bigr{)}\Bigr{)}\biggr{)}\Biggr{)},$
By substituting Eqs. A.4 and A into the energy-momentum tensor 4.6, we can easily derive the two structures, $T_{\mu\nu}T^{\mu\nu}$ and $T^{\mu}{}_{\mu}T^{\nu}{}_{\nu}$. The details of these structures are as follows:
$T_{\mu\nu}T^{\mu\nu}$ $=$ $\tfrac{1}{2}\bigl{(}\mathcal{P}\cosh(\gamma)+P_{1}\sinh(\gamma)\bigr{)}^{2}$ (A.6)
@@
$-8P_{1}\mathcal{P}(5479P_{1}^{2}+703\mathcal{P}^{2})\cosh(2\gamma)+16P_{1}\mathcal{P}(911P_{1}^{2}-125\mathcal{P}^{2})\cosh(4\gamma)-8P_{1}\mathcal{P}(281P_{1}^{2}+17\mathcal{P}^{2})\cosh(6\gamma)$
$+72P_{1}\mathcal{P}^{3}\cosh(8\gamma)+2(-2227P_{1}^{4}-306P_{1}^{2}\mathcal{P}^{2}+269\mathcal{P}^{4})\sinh(2\gamma)+2(1769P_{1}^{4}+4266P_{1}^{2}\mathcal{P}^{2}+253\mathcal{P}^{4})\sinh(4\gamma)$
@@
$-22P_{1}\mathcal{P}\cosh(4\gamma)+2(-53P_{1}^{2}+5\mathcal{P}^{2})\sinh(2\gamma)-11(P_{1}^{2}+\mathcal{P}^{2})\sinh(4\gamma)\bigr{)}\tanh(\gamma)\Bigr{)},$
@@
$+\lambda^{4}\Bigl{(}\tfrac{1}{36}d_{4}^{4}\bigl{(}\mathrm{csch}(\gamma)\bigr{)}^{6}\bigl{(}\mathrm{sech}(\gamma)\bigr{)}^{6}\bigl{(}\mathcal{P}\cosh(\gamma)+P_{1}\sinh(\gamma)\bigr{)}^{4}\bigl{(}-479P_{1}^{2}+113\mathcal{P}^{2}+140(4P_{1}^{2}+\mathcal{P}^{2})\cosh(2\gamma)$
@@
$+\frac{2d_{4}^{2}d_{7}\bigl{(}P_{1}+\mathcal{P}\coth(\gamma)\bigr{)}^{5}\bigl{(}-113P_{1}+43\mathcal{P}\coth(\gamma)\bigr{)}\bigl{(}\mathrm{sech}(\gamma)\bigr{)}^{3}}{9\bigl{(}\coth(\gamma)\bigr{)}^{2}}+\frac{3d_{11}d_{4}\bigl{(}P_{1}+\mathcal{P}\coth(\gamma)\bigr{)}^{6}}{\bigl{(}\coth(\gamma)\bigr{)}^{4}}$
$+\frac{16d_{7}^{2}\bigl{(}P_{1}+\mathcal{P}\coth(\gamma)\bigr{)}^{6}}{9\bigl{(}\coth(\gamma)\bigr{)}^{4}}\Bigr{)}.$

### v2

@@ Appendix A Details of the Lagrangian, and the Energy-Momentum Tensor
$\times\Bigl{(}3\mathcal{P}^{2}-13P_{1}^{2}+\cosh(4\gamma)(\mathcal{P}^{2}-3P_{1}^{2})+4\cosh(2\gamma)(\mathcal{P}^{2}+4P_{1}^{2})-16\bigl{(}\cosh(\gamma)\bigr{)}^{3}\mathcal{P}P_{1}\sinh(\gamma)\Bigr{)}$
$+\frac{4d_{4}^{4}}{\mathcal{P}}\biggl{(}6\mathrm{csch}(\gamma)(\mathcal{P}-P_{1})P_{1}^{2}(\mathcal{P}+P_{1})-5\bigl{(}\mathrm{csch}(\gamma)\bigr{)}^{7}\mathcal{P}^{4}$
$-\bigl{(}\mathrm{csch}(\gamma)\bigr{)}^{3}P_{1}^{2}(P_{1}^{2}-6\mathcal{P}^{2})+\bigl{(}\mathrm{csch}(\gamma)\bigr{)}^{5}(6\mathcal{P}^{2}P_{1}^{2}-4\mathcal{P}^{4})$
$+P_{1}^{2}\mathrm{sech}(\gamma)\Bigl{(}6(P_{1}^{2}-\mathcal{P}^{2})\tanh(\gamma)+(7P_{1}^{2}-12\mathcal{P}^{2})\bigl{(}\mathrm{sech}(\gamma)\bigr{)}^{2}\tanh(\gamma)$
$+4P_{1}\bigl{(}\mathrm{sech}(\gamma)\bigr{)}^{4}\bigl{(}8\mathcal{P}+5P_{1}\tanh(\gamma)\bigr{)}\Bigr{)}\biggr{)}\Biggr{)},$
By substituting Eqs. (A.4) and (A.5) into the energy-momentum tensor (4.6), we can easily derive the two structures, $T_{\mu\nu}T^{\mu\nu}$ and $T^{\mu}{}_{\mu}T^{\nu}{}_{\nu}$. The details of these structures are as follows:
$T_{\mu\nu}T^{\mu\nu}=\tfrac{1}{2}\bigl{(}\mathcal{P}\cosh(\gamma)+P_{1}\sinh(\gamma)\bigr{)}^{2}$
@@
$-8P_{1}\mathcal{P}(5479P_{1}^{2}+703\mathcal{P}^{2})\cosh(2\gamma)+16P_{1}\mathcal{P}(911P_{1}^{2}-125\mathcal{P}^{2})\cosh(4\gamma)$
$-8P_{1}\mathcal{P}(281P_{1}^{2}+17\mathcal{P}^{2})\cosh(6\gamma)+72P_{1}\mathcal{P}^{3}\cosh(8\gamma)$
$+2(-2227P_{1}^{4}-306P_{1}^{2}\mathcal{P}^{2}+269\mathcal{P}^{4})\sinh(2\gamma)+2(1769P_{1}^{4}+4266P_{1}^{2}\mathcal{P}^{2}+253\mathcal{P}^{4})\sinh(4\gamma)$
@@
$-22P_{1}\mathcal{P}\cosh(4\gamma)+2(-53P_{1}^{2}+5\mathcal{P}^{2})\sinh(2\gamma)-11(P_{1}^{2}+\mathcal{P}^{2})\sinh(4\gamma)\bigr{)}\tanh(\gamma)\Bigr{)},$ (A.6)
@@
$+\lambda^{4}\Bigl{(}\tfrac{1}{36}d_{4}^{4}\bigl{(}\mathrm{csch}(\gamma)\bigr{)}^{6}\bigl{(}\mathrm{sech}(\gamma)\bigr{)}^{6}\bigl{(}\mathcal{P}\cosh(\gamma)+P_{1}\sinh(\gamma)\bigr{)}^{4}$
$\times\bigl{(}-479P_{1}^{2}+113\mathcal{P}^{2}+140(4P_{1}^{2}+\mathcal{P}^{2})\cosh(2\gamma)$
@@
$+\frac{2d_{4}^{2}d_{7}\bigl{(}P_{1}+\mathcal{P}\coth(\gamma)\bigr{)}^{5}\bigl{(}-113P_{1}+43\mathcal{P}\coth(\gamma)\bigr{)}\bigl{(}\mathrm{sech}(\gamma)\bigr{)}^{3}}{9\bigl{(}\coth(\gamma)\bigr{)}^{2}}$
$+\frac{3d_{11}d_{4}\bigl{(}P_{1}+\mathcal{P}\coth(\gamma)\bigr{)}^{6}}{\bigl{(}\coth(\gamma)\bigr{)}^{4}}+\frac{16d_{7}^{2}\bigl{(}P_{1}+\mathcal{P}\coth(\gamma)\bigr{)}^{6}}{9\bigl{(}\coth(\gamma)\bigr{)}^{4}}\Bigr{)}.$